# Massive Higher-Spin Fields in the Fractional Quantum Hall Effect

### v1

@@ Appendix B Non-relativistic limit
When writing the terms in the action (34) in complex form the following identities will be useful
$Y_{a_{1}\cdots a_{k}}Z_{a_{1}\cdots a_{k}}=2^{k-2}\left(\mathcal{Y}_{[k]}^{*}\mathcal{Z}_{[k]}+\mathcal{Y}_{[k]}\mathcal{Z}_{[k]}^{*}\right),$ (46)
$Y_{b_{1}\cdots b_{k}}\partial_{a}Z_{ab_{1}\cdots b_{k}}=2^{k-2}\left(\mathcal{Y}_{[k]}^{*}\bar{\partial}\mathcal{Z}_{[k+1]}+\mathcal{Y}_{[k]}\partial\mathcal{Z}_{[k+1]}^{*}\right),$
$\partial_{a}Z_{ac_{1}\cdots c_{k-1}}\partial_{b}Z_{bc_{1}\cdots c_{k-1}}=\frac{1}{2}\partial_{a}Z_{b_{1}\cdots b_{k}}\partial_{a}Z_{b_{1}\cdots b_{k}}=2^{k-1}\bar{\partial}\mathcal{Z}_{[k]}\partial\mathcal{Z}_{[k]}^{*},$
where in the last identity we have neglected total derivative terms and defined the complex derivatives $\partial=\partial_{1}+i\partial_{2}$ and ¯$\bar{\partial}=\partial_{1}-i\partial_{2}$ and ${\rm T.D.}$ stands for total derivative terms that we neglect.
Using the prescription (45), we define complex fields $\mathcal{S}_{[i]}$ and $\mathcal{V}_{[j]}$ associated to the fields $S_{a_{1}\cdots b_{i}}$ and $v_{a_{1}\cdots a_{j}}$ entering in the parametrisation (42), (43) and (44) of the relativistic fields. The non-relativistic limit will be carried out after introducing the following field redefinition for the fields (45)
$\mathcal{S}_{[i]}=e^{-i(mc^{2}-E_{0})t}\mathcal{Q}_{[i]},\qquad\mathcal{V}_{[j]}=e^{-i(mc^{2}-E_{0})t}\mathcal{U}_{[j]},$ (47)
where $m$ coincides with the mass parameter in the action (34) and $E_{0}$ is an arbitrary constant. This, in turn, introduces new vectors and symmetric traceless tensors $Q_{a_{1}\cdots a_{i}}$ and $u_{a_{1}\cdots a_{j}}$, whose components are given by the real and imaginary parts of $\mathcal{Q}_{[k]}$ and $\mathcal{U}_{[k]}$ as indicated in (45). In the following we will show that when considering the Lagrangian for a spin-2 (spin-4) field and taking the non-relativistic limit defined by implementing (47) and setting $c\rightarrow\infty$, the field $Q_{ab}$ ($Q_{abcd}$) decouples from all the lower-spin fields and is described by the a Schroedinger action of the form given in Eq. (17).
A Spin-2 case
We start considering the Lagrangian (34) for a massive traceless spin-2 field. Using the $\lambda$-symmetry to gauge the scalar field $S$ away we find the action
$\kappa^{-1}\mathcal{L}=\frac{1}{2}h^{\mu\nu}\Box h_{\mu\nu}-h_{\mu\lambda}\partial^{\mu}\partial_{\nu}h^{\nu\lambda}-\frac{m^{2}c^{2}}{2}h^{\mu\nu}h_{\mu\nu}+\frac{1}{2}B_{\mu}\Box B^{\mu}-\frac{1}{2}B^{\mu}\partial_{\mu}\partial_{\nu}B^{\nu}+\sqrt{2}mc\;h^{\mu\nu}\partial_{\mu}B_{\nu}.$ (48)
We adopt the parametrisation (42) and use the complex notation (45). Thus, the identities (46) allow to write the action as (up to total derivatives)
$\kappa^{-1}\mathcal{L}$ $=\frac{3}{4c^{6}}\dot{\phi}^{2}-\frac{3}{4c^{4}}\partial\phi\bar{\partial}\phi+\frac{1}{c^{2}}\dot{\mathcal{S}}_{[2]}\dot{\mathcal{S}}_{[2]}^{*}-\frac{1}{c^{6}}\dot{\phi}^{2}+\frac{1}{4c^{2}}\partial\phi\bar{\partial}\phi+\frac{1}{2c^{2}}\left(\bar{\partial}\mathcal{S}_{[2]}\bar{\partial}\phi+\partial\mathcal{S}_{[2]}^{*}\partial\phi\right)$ (49)
$-\frac{3m^{2}}{4c^{2}}\phi^{2}-m^{2}c^{2}\mathcal{S}_{[2]}\mathcal{S}^{*}_{[2]}+\frac{1}{2c^{2}}\left(\dot{\mathcal{S}}_{[1]}\dot{\mathcal{S}}_{[1]}^{*}-\dot{\mathcal{S}}_{[1]}\bar{\partial}v-\dot{\mathcal{S}}_{[1]}^{*}\partial v+\partial v\bar{\partial}v\right)$
$-\frac{1}{4}\left(\partial\mathcal{S}_{[1]}\bar{\partial}\mathcal{S}_{[1]}^{*}+\bar{\partial}\mathcal{S}_{[1]}\partial\mathcal{S}_{[1]}^{*}\right)+\frac{1}{8}\left(\partial\mathcal{S}^{*}_{[1]}+\bar{\partial}\mathcal{S}_{[1]}\right)^{2}+\frac{\sqrt{2}m}{c^{3}}\phi\dot{v}$
$+\frac{m}{2\sqrt{2}c}\phi\left(\partial\mathcal{S}^{*}_{[1]}+\bar{\partial}\mathcal{S}_{[1]}\right)-\frac{mc}{\sqrt{2}}\left(\bar{\partial}\mathcal{S}_{[2]}\mathcal{S}_{[1]}^{*}+\partial\mathcal{S}_{[2]}^{*}\mathcal{S}_{[1]}\right)+\kappa^{-1}\mathcal{L}_{\mathcal{V}_{[1]}},$
where we have isolated the terms that depend on $\mathcal{V}_{[1]}$ in the Lagrangian
$\kappa^{-1}\mathcal{L}_{\mathcal{V}_{[1]}}$ $=m^{2}\mathcal{V}_{[1]}\mathcal{V}^{*}_{[1]}+\frac{1}{2c^{2}}\left(\partial\mathcal{V}_{[1]}\bar{\partial}\mathcal{V}_{[1]}^{*}+\bar{\partial}\mathcal{V}_{[1]}\partial\mathcal{V}_{[1]}^{*}\right)-\frac{1}{c^{2}}\left(\bar{\partial}\mathcal{S}_{[2]}\dot{\mathcal{V}}_{[1]}^{*}+\partial\mathcal{S}_{[2]}^{*}\dot{\mathcal{V}}_{[1]}\right)$ (50)
$+\frac{1}{2c^{4}}\left(\dot{\mathcal{V}}_{[1]}\bar{\partial}\phi+\dot{\mathcal{V}}^{*}_{[1]}\partial\phi\right)-\frac{1}{4c^{2}}\left(\partial\mathcal{V}^{*}_{[1]}+\bar{\partial}\mathcal{V}_{[1]}\right)^{2}-\frac{m}{\sqrt{2}c}\left(\mathcal{V}_{[1]}\bar{\partial}v+\mathcal{V}_{[1]}^{*}\partial v+\dot{\mathcal{S}}_{[1]}\mathcal{V}_{[1]}^{*}+\dot{\mathcal{S}}_{[1]}^{*}\mathcal{V}_{[1]}\right).$
Now we introduce complex fields in terms of $\mathcal{Q}_{[2]}$, $\mathcal{Q}_{[1]}$ and $\mathcal{U}_{[1]}$ by means of the field redefinition (47), in terms of which the Lagrangian takes the form
$\kappa^{-1}\mathcal{L}$ $=-2mE_{0}\mathcal{Q}_{[2]}\mathcal{Q}_{[2]}^{*}-2im\mathcal{Q}_{[2]}\dot{\mathcal{Q}}_{[2]}^{*}-mE_{0}\mathcal{Q}_{[1]}\mathcal{Q}_{[1]}^{*}-im\mathcal{Q}_{[1]}\dot{\mathcal{Q}}_{[1]}^{*}$ (51)
$+\frac{m^{2}c^{2}}{2}\mathcal{Q}_{[1]}\mathcal{Q}_{[1]}^{*}-\frac{mc}{\sqrt{2}}\left(\bar{\partial}\mathcal{Q}_{[2]}\mathcal{Q}_{[1]}^{*}+\partial\mathcal{Q}_{[2]}^{*}\mathcal{Q}_{[1]}\right)+\kappa^{-1}\mathcal{L}_{\mathcal{U}_{[1]}}$
$-\frac{1}{4}\left(\partial\mathcal{Q}_{[1]}\bar{\partial}\mathcal{Q}_{[1]}^{*}+\bar{\partial}\mathcal{Q}_{[1]}\partial\mathcal{Q}_{[1]}^{*}\right)+\frac{1}{8}\left(e^{i(mc^{2}-E_{0})t}\partial\mathcal{Q}^{*}_{[1]}+e^{-i(mc^{2}-E_{0})t}\bar{\partial}\mathcal{Q}_{[1]}\right)^{2}+O(1/c).$
Note that the field redefinition (47) introduced some factors $e^{-i(mc^{2}-E_{0})t}$ that will oscillate in the $c\to\infty$ limit. On the other hand, these terms will decouple in non-relativistic the limit, so that we will ignore this issue as in [53]. After introducing the new variables, the Lagrangian for $\mathcal{U}_{[1]}$ now has the form
$\kappa^{-1}\mathcal{L}_{\mathcal{U}_{[1]}}$ $=m^{2}\mathcal{U}_{[1]}\mathcal{U}^{*}_{[1]}-\mathcal{U}_{[1]}\mathcal{J}_{[1]}^{*}-\mathcal{U}_{[1]}^{*}\mathcal{J}_{[1]}+O(1/c^{2}),$ (52)
where we have defined
$\mathcal{J}_{[1]}$ $=-\frac{im^{2}c}{\sqrt{2}}\mathcal{Q}_{[1]}+im\bar{\partial}\mathcal{Q}_{[2]}+\frac{1}{c}\left(\frac{m}{\sqrt{2}}\partial v+\frac{imE_{0}}{\sqrt{2}}\mathcal{Q}_{[1]}+\frac{m}{\sqrt{2}}\dot{\mathcal{Q}}_{[1]}\right)+O(1/c^{2}),$ (53)
Now we note that for large values of $c$ the field equations for $\mathcal{U}_{[1]}^{*}$ lead to
$\mathcal{U}_{[1]}=\frac{1}{m^{2}}\mathcal{J}_{[1]}+O(1/c^{2}),$ (54)
and therefore integrating out $\mathcal{U}_{[1]}$ yields the following effective Lagrangian
$\kappa^{-1}\mathcal{L}^{eff}_{\mathcal{U}_{[1]}}$ $=-\frac{1}{m^{2}}\mathcal{J}_{[1]}\mathcal{J}^{*}_{[1]}=-\frac{m^{2}c^{2}}{2}\mathcal{Q}_{[1]}\mathcal{Q}_{[1]}^{*}+\frac{mc}{\sqrt{2}}\left(\mathcal{Q}_{[1]}\partial\mathcal{Q}_{[2]}^{*}+\mathcal{Q}_{[1]}^{*}\bar{\partial}\mathcal{Q}_{[2]}\right)-\bar{\partial}\mathcal{Q}_{[2]}\partial\mathcal{Q}_{[2]}^{*}$ (55)
$+\frac{im}{2}\mathcal{Q}_{[1]}\left(\bar{\partial}v+\dot{\mathcal{Q}}_{[1]}^{*}-iE_{0}\mathcal{Q}_{[1]}^{*}\right)-\frac{im}{2}\mathcal{Q}_{[1]}^{*}\left(\partial v+\dot{\mathcal{Q}}_{[1]}+iE_{0}\mathcal{Q}_{[1]}\right)+O(1/c),$
When replacing this expression by $\mathcal{L}_{\mathcal{U}_{[1]}}$ in the Lagrangian (51), the divergent terms cancel and $\mathcal{Q}_{[2]}$ decouples from the other fields. Thus, in the limit $c\to\infty$, up to total derivative terms, we find the spin-2 Schroedinger Lagrangian Eq. (26)
$\mathcal{L}=2i\kappa m\dot{\mathcal{Q}}_{[2]}\mathcal{Q}_{[2]}^{*}-\kappa\bar{\partial}\mathcal{Q}_{[2]}\partial\mathcal{Q}_{[2]}^{*}-2\kappa mE_{0}\mathcal{Q}_{[2]}\mathcal{Q}_{[2]}^{*}+\ldots$ (56)
where the dots denote terms that do not depend on $\mathcal{Q}_{[2]}$. Using the fact that up to total derivatives $\epsilon_{bc}Q_{ab}\dot{Q}_{ac}=-2i\dot{\mathcal{Q}}_{[2]}\mathcal{Q}^{*}_{[2]}$ and the identities (46), the $\mathcal{Q}_{[2]}$-dependent part of this Lagrangian can be put in the form (4) with coefficients $k_{1}=-8\kappa m$, $k_{2}=4\kappa$ and $k_{3}=2\kappa mE_{0}$.
B Spin-4 case
The Lagrangian (34) evaluated for $D=2+1$ and $s=4$ can be written as
$\mathcal{L}=\kappa\sum_{k=0}^{4}\left(K_{\varphi_{\mu(k)}}+M_{\varphi_{\mu(k)}\varphi_{\nu(k-1)}}\right),$ (57)
where, using (41), the different kinetic and mixing terms different terms take the form
$K_{\Phi}=\frac{1}{2}\Phi^{\mu\nu\rho\sigma}\Box\Phi_{\mu\nu\rho\sigma}-2\Phi^{\mu\nu\rho\sigma}\partial_{\mu}\partial^{\lambda}\Phi_{\lambda\nu\rho\sigma}-m^{2}c^{2}\,\Phi^{\mu\nu\rho\sigma}\Phi_{\mu\nu\rho\sigma},$ (58)
$K_{\chi}=\frac{1}{2}\chi^{\mu\nu\rho}\Box\chi_{\mu\nu\rho}-\frac{3}{2}\chi^{\mu\nu\rho}\partial_{\mu}\partial^{\lambda}\chi_{\lambda\nu\rho},$
$K_{h}=\frac{1}{2}h^{\mu\nu}\Box h_{\mu\nu}-h^{\mu\nu}\partial_{\mu}\partial^{\lambda}h_{\lambda\nu}+\frac{7}{10}\,m^{2}c^{2}\,h^{\mu\nu}h_{\mu\nu},$
$K_{B}=\frac{1}{2}B^{\mu}\Box B_{\mu}-\frac{1}{2}B^{\mu}\partial_{\mu}\partial^{\lambda}B_{\lambda}+2\,m^{2}c^{2}\,B^{\mu}B_{\mu},$
$K_{S}=\frac{1}{2}S\Box S+\frac{15}{2}\,m^{2}c^{2}\,S^{2},$
and
$M_{\Phi\chi}$ $=2\,mc\,\Phi^{\mu\nu\rho\sigma}\partial_{\mu}\chi_{\nu\rho\sigma},$ $M_{\chi h}$ $=\frac{6}{\sqrt{5}}\,mc\,\chi^{\mu\nu\rho}\partial_{\mu}h_{\nu\rho},$ (59)
$M_{hB}$ $=\sqrt{10}\,mc\,h^{\mu\nu}\partial_{\mu}B_{\nu},$ $M_{BS}$ $=2\sqrt{2}\,mc\,B^{\mu}\partial_{\mu}S.$
Using the parametrisation (42), (43), (44), together with the identities (46) we find that in complex notation the kinetic terms and the mixing terms can be written in a unified notation for $k\geq 2$ as
$K_{\varphi_{\mu(k)}}=\frac{2^{k-1}}{2c^{2}}\dot{\mathcal{S}}_{[k]}\dot{\mathcal{S}}_{[k]}^{*}-\frac{2^{k-3}k}{c^{2}}\left(\dot{\mathcal{V}}_{[k-1]}\partial\mathcal{S}_{[k]}^{*}+\dot{\mathcal{V}}_{[k-1]}^{*}\bar{\partial}\mathcal{S}_{[k]}\right)+2^{k-2}\left(\frac{k}{2}-1\right)\partial_{a}\mathcal{S}_{[k]}\partial_{a}\mathcal{S}_{[k]}^{*}$ (60)
$+2^{k-3}\,k\;m_{k}\,m^{2}\,\mathcal{V}_{[k-1]}\mathcal{V}_{[k-1]}^{*}-2^{k-2}m_{k}\,m^{2}c^{2}\,\mathcal{S}_{[k]}\mathcal{S}_{[k]}^{*}+O(1/c^{2}),$
$M_{\varphi_{\mu(k)}\varphi_{\nu(k-1)}}=-2^{k-3}c_{k}\left(\frac{m}{c}\mathcal{V}_{[k-1]}^{*}\dot{\mathcal{S}}_{[k-1]}+\frac{m}{c}\mathcal{V}_{[k-1]}\dot{\mathcal{S}}_{[k-1]}^{*}+mc\,\mathcal{S}_{[k-1]}\partial\mathcal{S}_{[k]}^{*}+mc\,\mathcal{S}_{[k-1]}^{*}\bar{\partial}\mathcal{S}_{[k]}\right)+O(1/c^{2}).$
where we have reinstated the coefficients $m_{k}$ and $c_{k}$, whose explicit form is given in Eq. (35). The non-relativistic limit can be taken after redefining the variables by means of (47), in terms of which (60) takes the form
$K_{\varphi_{\mu(k)}}=2^{k-1}im\dot{\mathcal{Q}}_{[k]}\mathcal{Q}_{[k]}^{*}+2^{k-1}\left(\frac{m^{2}c^{2}}{2}-mE_{0}\right)\mathcal{Q}_{[k]}\mathcal{Q}_{[k]}^{*}+2^{k-3}im\,k\Big{(}\mathcal{U}_{[k-1]}\partial\mathcal{Q}_{[k]}^{*}-\mathcal{U}_{[k-1]}^{*}\bar{\partial}\mathcal{Q}_{[k]}\Big{)}$ (61)
$+2^{k-2}\left(\frac{k}{2}-1\right)\partial_{a}\mathcal{Q}_{[k]}\partial_{a}\mathcal{Q}_{[k]}^{*}+2^{k-3}\,k\;m_{k}\,m^{2}\,\mathcal{U}_{[k-1]}\mathcal{U}_{[k-1]}^{*}-2^{k-2}m_{k}\,m^{2}c^{2}\,\mathcal{Q}_{[k]}\mathcal{Q}_{[k]}^{*}+O(1/c^{2}),$
$M_{\varphi_{\mu(k)}\varphi_{\nu(k-1)}}=-2^{k-3}c_{k}\,mc\,\bigg{(}-im\left(\mathcal{U}_{[k-1]}^{*}\mathcal{Q}_{[k-1]}-\mathcal{U}_{[k-1]}\mathcal{Q}_{[k-1]}^{*}\right)+\mathcal{Q}_{[k-1]}\partial\mathcal{Q}_{[k]}^{*}+\mathcal{Q}_{[k-1]}^{*}\bar{\partial}\mathcal{Q}_{[k]}\bigg{)}+O(1/c^{2}).$
@@ B Spin-4 case
From these expressions we see that in order to eliminate the divergent term contained in $K_{\varphi_{\mu(s)}}$ term, the constants $m_{s}$ should take precisely the form (35). Moreover, the term $M_{\varphi_{\mu(s)}\varphi_{\mu(s-1)}}$ also contains divergent terms, which disappear after integrating out $\mathcal{U}_{[s-1]}$ from the action (57). An inspection of the different terms in (61) shows that the terms in the action involving the $\mathcal{U}_{[s-1]}$ have the following structure
$\mathcal{L}_{\mathcal{U}_{[s-1]}}=2^{s-3}\kappa\bigg{[}s\;m_{s}\,m^{2}\,\mathcal{U}_{[s-1]}\mathcal{U}_{[s-1]}^{*}-im\mathcal{U}_{[s-1]}\bigg{(}c_{s}\,mc\,\mathcal{Q}_{[s-1]}^{*}-s\partial\mathcal{Q}_{[s]}^{*}\bigg{)}+im\mathcal{U}_{[s-1]}^{*}\bigg{(}c_{s}\,mc\,\mathcal{Q}_{[s-1]}-s\bar{\partial}\mathcal{Q}_{[s]}\bigg{)}\bigg{]}+O(1/c^{2})$ (62)
@@
from which we find the following field equation of $\mathcal{U}_{[s-1]}$ for large values of $c$
$\mathcal{U}_{[s-1]}=-\frac{i}{s\;m_{s}}\bigg{(}c_{s}\,c\mathcal{Q}_{[s-1]}-\frac{s}{m}\bar{\partial}\mathcal{Q}_{[s]}\bigg{)}+O(1/c).$ (63)
@@
$\mathcal{L}^{eff}_{\mathcal{U}_{[s-1]}}=-\frac{2^{s-3}\kappa}{s\;m_{s}}\bigg{(}c_{s}\,mc\,\mathcal{Q}_{[s-1]}-s\bar{\partial}\mathcal{Q}_{[s]}\bigg{)}\bigg{(}c_{s}\,mc\,\mathcal{Q}_{[s-1]}^{*}-s\partial\mathcal{Q}_{[s]}^{*}\bigg{)}+\cdots,$ (64)
where $\cdots$ represents either terms that do not involve $\mathcal{Q}_{[s]}$ or are subleading in $1/c$. Replacing (62) by (64), inserting it back in the action (57), and taking the limit $c\rightarrow\infty$ gives rise to the following action for $\mathcal{Q}_{[s]}$
$\mathcal{L}=2^{s-1}\kappa\;m\bigg{[}i\dot{\mathcal{Q}}_{[s]}\mathcal{Q}_{[s]}^{*}-\frac{1}{2m}\bar{\partial}\mathcal{Q}_{[s]}\partial\mathcal{Q}_{[s]}^{*}-E_{0}\mathcal{Q}_{[s]}\mathcal{Q}_{[s]}^{*}\bigg{]}+\cdots,$ (65)
where $\mathcal{Q}_{[s]}$ is decoupled from all the other fields in the Lagrangian. Even though this result was considered for $s=4$, it is written in a general form that allows to generalize the result to arbitrary values of $s$. Using the identities (46) plus the fact that up to total derivatives $\epsilon_{mn}Q_{abcm}\dot{Q}_{abcn}=-8i\dot{\mathcal{Q}}_{[4]}\mathcal{Q}^{*}_{[4]}$, the Lagrangian for $\mathcal{Q}_{[4]}$ takes the form (8) with parameters $k^{\prime}_{1}=-8\kappa m$, $k^{\prime}_{2}=4\kappa$ and $k^{\prime}_{3}=2\kappa mE_{0}$.

### v2

@@ Appendix C Non-relativistic limit of the Proca-like actions
Using the prescription (61), we then define complex fields $\mathcal{S}_{[k]}$ and $\mathcal{V}_{[k-1]}$ associated to the fields $S_{a_{1}\cdots b_{k}}$ and $v_{a_{1}\cdots a_{k-1}}$ denoting the two spatial components of highest rank entering the decomposition of each field in the Lagrangian (46). Similar complex fields can be introduced also for the components with less spatial indices (with the exception of the scalar ones). In analogy with what we discussed in the spin-2 case, the latter however always decouple in the limit $c\to\infty$. As for spin two, before taking the non-relativistic limit we then introduce the field redefinition (28) for all fields $\mathcal{S}_{[k]}$ and $\mathcal{V}_{[k-1]}$. This, in turn, introduces new symmetric traceless tensors $Q_{a_{1}\cdots a_{i}}$ and $u_{a_{1}\cdots a_{j}}$, whose components are given by the real and imaginary parts of $\mathcal{Q}_{[k]}$ and $\mathcal{U}_{[k]}$ as indicated in (61).
After these redefinitions, the Lagrangian (46), evaluated for $D=2+1$, becomes
$\kappa^{-1}\mathcal{L}$ $=\sum_{k=1}^{s}2^{k-2}\bigg{[}im\dot{\mathcal{Q}}_{[k]}\mathcal{Q}_{[k]}^{*}+\left(\frac{m^{2}c^{2}}{2}(1-m_{k})-mE_{0}\right)|\mathcal{Q}_{[k]}|^{2}+\frac{1}{2}\left(\frac{k}{2}-1\right)|\partial\mathcal{Q}_{[k]}|^{2}$ (62)
$+\frac{k}{2}\;m_{k}\,m^{2}|\mathcal{U}_{[k-1]}|^{2}+\frac{im\,k}{2}\mathcal{U}_{[k-1]}\partial\mathcal{Q}_{[k]}^{*}+\frac{c_{k}mc}{2}\bigg{(}im\,\mathcal{U}_{[k-1]}^{*}\mathcal{Q}_{[k-1]}-\mathcal{Q}_{[k-1]}\partial\mathcal{Q}_{[k]}^{*}\bigg{)}\bigg{]}+\rm{c.c.}+O(c^{-2})$
@@
Notice that in order to eliminate the term diverging like $c^{2}$ in the contribution with $k=s$ in Eq. (62), the constant $m_{s}$ should take precisely the form (47a). The contribution with $k=s$ contains another divergent term linear in $c$, which however cancels after integrating out $\mathcal{U}_{[s-1]}$ from the action. Indeed, inspection of Eq. (62) shows that the portion of the Lagrangian involving $\mathcal{U}_{[s-1]}$ has the following structure:
$\frac{2^{3-s}}{\kappa}\mathcal{L}_{\mathcal{U}_{[s-1]}}$ $=im\mathcal{U}_{[s-1]}^{*}\bigg{(}c_{s}\,mc\,\mathcal{Q}_{[s-1]}-s\bar{\partial}\mathcal{Q}_{[s]}\bigg{)}$ (63)
$+\frac{s\,m_{s}m^{2}}{2}|\mathcal{U}_{[s-1]}|^{2}+\rm{c.c.}+O(c^{-2})$
@@
from which we find the following field equation for $\mathcal{U}_{[s-1]}$ for large values of $c$:
$\mathcal{U}_{[s-1]}=-\frac{i}{s\;m_{s}}\bigg{(}c_{s}\,c\mathcal{Q}_{[s-1]}-\frac{s}{m}\bar{\partial}\mathcal{Q}_{[s]}\bigg{)}+O(c^{-1}).$ (64)
@@
$\mathcal{L}^{eff}_{\mathcal{U}_{[s-1]}}=-\frac{2^{s-3}\kappa}{s\;m_{s}}\bigg{|}c_{s}\,mc\,\mathcal{Q}_{[s-1]}-s\bar{\partial}\mathcal{Q}_{[s]}\bigg{|}^{2}+\cdots,$ (65)
where $\cdots$ represents either terms that do not involve $\mathcal{Q}_{[s]}$ or are subleading in $1/c$. Replacing (63) by (65) in the Lagrangian (62), then cancels the term diverging linearly in $c$ coming from the contribution with $k=s$. A similar mechanism allows to cancel all divergences in $c$ by integrating out the fields $\mathcal{U}_{[k]}$ recursively. Eventually, taking the limit $c\rightarrow\infty$ gives
$\mathcal{L}=2^{s-1}\kappa\;m\bigg{[}i\dot{\mathcal{Q}}_{[s]}\mathcal{Q}_{[s]}^{*}-\frac{1}{2m}|\partial\mathcal{Q}_{[s]}|^{2}-E_{0}|\mathcal{Q}_{[s]}|^{2}\bigg{]}+\dots,$ (66)
where $\mathcal{Q}_{[s]}$ is decoupled from all the other fields in the Lagrangian.
Even though we focused on the $s=4$ case, the previous derivation is written in a general form that allows one to generalize the result to arbitrary values of $s$. Moreover, as anticipated in the main text, using, e.g., that up to total derivatives $\epsilon_{mn}Q_{abcm}\dot{Q}_{abcn}=-8i\dot{\mathcal{Q}}_{[4]}\mathcal{Q}^{*}_{[4]}$, the Lagrangian for $\mathcal{Q}_{[4]}$ takes the form in Eq. (8) with parameters $k^{\prime}_{1}=-8\kappa m$, $k^{\prime}_{2}=4\kappa$ and $k^{\prime}_{3}=2\kappa mE_{0}$.